# Rethinking Disentanglement under Dependent Factors of Variation

### v2

@@ Abstract
Representation learning enables the discovery and extraction of underlying factors of variation from data. A representation is typically considered disentangled when it isolates these factors in a way that is interpretable to humans. Existing definitions and metrics for disentanglement often assume that the factors of variation are statistically independent. However, this assumption rarely holds in real-world settings, limiting the applicability of such definitions and metrics to ideal scenarios. In this work, we propose a novel definition of disentanglement grounded in information theory, which remains valid even when the factors are dependent. We show that this definition is equivalent to requiring the representation to consist of minimal and sufficient variables. Based on this formulation, we introduce a method to quantify the degree of disentanglement that remains effective in the presence of statistical dependencies among factors. Through a series of experiments, we demonstrate that our method reliably measures disentanglement in both independent and dependent settings, where existing approaches fail under the latter.
@@ 1 Introduction
In representation learning, data are commonly assumed to be characterized by a set of factors of variation (henceforth, factors) and nuisances. The distinction between them lies in task relevance: factors directly influence the outcome of interest, whereas nuisances do not. For instance, in a dataset of fruit images where the task is to describe the fruits, relevant factors may include fruit type, color, or size, while irrelevant nuisances might be the background color or the shadows cast. A large body of work has emphasized the importance of learning disentangled representations, where distinct dimensions of the representation correspond to different factors of variation (Schmidhuber, 1992; Bengio et al., 2013; Ridgeway, 2016; Lake et al., 2017; Achille & Soatto, 2018a). Such representations are valuable across a range of applications, including: (i) interpretability of learned features and model predictions (Hsu et al., 2017; Worrall et al., 2017; Gilpin et al., 2018; Zhang & Zhu, 2018; Liu et al., 2020; Zhu et al., 2021; Nauta et al., 2023; Almudévar et al., 2024), (ii) fairer decision-making by mitigating or removing biases in sensitive attributes such as gender or race (Calders & Žliobaitė, 2013; Barocas & Selbst, 2016; Kumar et al., 2017; Locatello et al., 2019a; Sarhan et al., 2020; Chen et al., 2023), and (iii) empowering generative models to synthesize data with controllable attributes (Yan et al., 2016; Paige et al., 2017; Huang et al., 2018; Dupont, 2018; Kazemi et al., 2019; Antoran & Miguel, 2019; Zhou et al., 2020; Shen et al., 2022; Almudévar et al., 2024).
The definition in the previous paragraph provides an intuitive, though imprecise, understanding of disentangled representations. It raises two fundamental questions: (i) what does it precisely mean to separate the factors, and (ii) how can we evaluate whether the factors are indeed separated? In Section 2, we review several works that attempt to address these two questions. Most of these approaches assume that the factors of variation are independent—both from one another and from the nuisances. Consequently, the corresponding definitions and evaluation metrics are grounded in this assumption of independence. Many of the benchmark datasets used to validate these methods also exhibit (approximately) independent factors (LeCun et al., 2004; Liu et al., 2015; Reed et al., 2015; Matthey et al., 2017; Burgess & Kim, 2018; Gondal et al., 2019), thereby making the independence-based definitions and metrics largely applicable in those contexts.
@@
  • We prove that these properties hold if and only if the representations are both minimal and sufficient. Moreover, we argue that evaluating minimality and sufficiency provides a more meaningful assessment of disentanglement than analyzing the properties in isolation.
  • We propose a method to quantify the degree of minimality and sufficiency, enabling the evaluation of disentanglement in arbitrary representations.  [^1] https://github.com/antonioalmudevar/dependent_disentanglement
Figure 1: If the factors are dependent, then—even with a perfectly disentangled encoder—each sub-representation will still contain information about other factors. For example, if $z_{1}$ encodes the shape of a banana, the fruit’s color is most likely yellow, possibly green if unripe, and almost never red. Likewise, if $z_{2}$ encodes the color yellow, the shape is most likely that of a banana or a lemon, but very unlikely to be that of a strawberry.
@@ 2 Related Work
Despite being a topic of great interest, there is no widely agreed-upon definition of disentangled representations. Intuitively, a disentangled representation separates the different factors of variation underlying the data (Desjardins et al., 2012; Bengio et al., 2013; Cohen & Welling, 2014; Kulkarni et al., 2015; Ridgeway, 2016). Early evaluations of disentanglement were largely based on visual inspection (Kingma & Welling, 2013). However, more precise definitions are necessary to understand the limitations and practical relevance of disentangled representations in downstream applications. One of the earliest and most widely cited definitions is given in Bengio et al. (2013), and has been adopted by several subsequent works (Higgins et al., 2017; Kim & Mnih, 2018; Kim et al., 2019; Suter et al., 2019). This definition states that a disentangled representation is one in which a change in a single dimension of the representation corresponds to a change in a single factor of variation, while remaining invariant to changes in other factors. An alternative view argues that disentanglement should be defined such that a change in a single factor of variation leads to a change in only one dimension of the representation (Kumar et al., 2017; Locatello et al., 2019b). More recent efforts define disentanglement based on a set of desirable properties. Notably, Ridgeway & Mozer (2018) and Eastwood & Williams (2018) independently proposed three key properties that a representation should satisfy to be considered disentangled. While the terminology differs slightly between the two works, the underlying ideas align: (i) Modularity(referred to as disentanglement in Eastwood & Williams (2018)): each representation variable captures at most one factor of variation; (ii) Compactness(completeness in Eastwood & Williams (2018)): each factor of variation is captured by a single representation variable; and (iii) Explicitness(informativeness in Eastwood & Williams (2018)): the representation retains all task-relevant information about the factors present in the input. All of the above definitions present two key limitations: (i) they do not account for invariance to nuisance factors that may be present in the input; and (ii) they assume that the factors of variation are independent of one another and of the nuisances—which generally does not hold in real-world data. We propose a set of properties that a representation must satisfy to be considered disentangled, explicitly addressing both of these aspects.
@@
  • Intervention-based Metrics. These methods construct subsets of data samples that share a common factor of variation while differing in others. Representations for these subsets are then computed and compared to produce a disentanglement score. Representative methods in this category include $\beta$-VAE (Higgins et al., 2017), FactorVAE (Kim & Mnih, 2018), and R-FactorVAE (Kim et al., 2019). These metrics primarily aim to assess modularity but do not account for compactness or explicitness (Sepliarskaia et al., 2019).
  • Information-based Metrics. These approaches estimate mutual information between the factors and individual variables in the representation. Widely used examples include the Mutual Information Gap (MIG) (Chen et al., 2018) and Robust MIG (Do & Tran, 2019). These metrics are mainly designed to capture compactness.
  • Predictor-based Metrics. These methods train regressors or classifiers to predict the factors from the learned representations. The resulting predictors are then analyzed to determine the contribution of each representation variable to each factor. Separated Attribute Predictability (SAP) (Kumar et al., 2017), focuses exclusively on compactness. Subsequently, more comprehensive frameworks have been proposed to evaluate multiple properties of disentanglement. Notably, Eastwood & Williams (2018) introduced DCI (Disentanglement, Completeness, and Informativeness), while Ridgeway & Mozer (2018) proposed an equivalent set: Modularity, Compactness, and Explicitness.
@@
The Information Bottleneck (IB) framework (Tishby et al., 2000) formalizes a trade-off between fidelity and compression in learned representations. A representation is deemed sufficient if it retains all information relevant to the target task, and minimal if it excludes all information irrelevant to it. Formally, given an input $x$, target variable $y$, and representation $z$, sufficiency corresponds to $I(z;y)=I(x;y)$, while minimality requires $I(z;x)=I(z;y)$. The IB principle has been leveraged in a range of works to design regularizers for representation learning (Alemi et al., 2016; 2018; Achille & Soatto, 2018b; a; Almudévar et al., 2025), and has also been applied to promote disentanglement (Vera et al., 2018; Yamada et al., 2020; Jeon et al., 2021; Gao et al., 2021). To the best of our knowledge, however, this is the first work to formally define disentanglement in terms of minimal sufficient representations. In our formulation, each factor of variation is treated as a distinct target variable, and the overall representation is decomposed into sub-representations. We define a representation as disentangled if each sub-representation is both minimal and sufficient with respect to its corresponding factor.
@@ 3 Desirable Properties of a Disentangled Representation
In our representation learning setup, we consider raw data $\bm{x}$, which is fully explained by a set of task-relevant factors $\bm{y}=\{y_{i}\}_{i=1}^{n}$ and a set of nuisances $\bm{n}$. The factors $\bm{y}$ correspond to the underlying sources or causes that influence the observed data and are relevant to the task at hand. We assume that these factors are conditionally independent given $\bm{x}$, i.e., $p(y_{i},y_{j}\mid\bm{x})=p(y_{i}\mid\bm{x})p(y_{j}\mid\bm{x})$ for $i\neq j$ [^2] Most prior work assumes marginal independence between factors, i.e., $p(y_{i},y_{j})=p(y_{i})p(y_{j})$ for $i\neq j$, which rarely holds in real-world scenarios.. The nuisances $\bm{n}$ refer to sources of variation in $\bm{x}$ that are irrelevant to the task.
We define a representation $\bm{z}$ as a random variable governed by the conditional distribution $p(\bm{z}\mid\bm{x})$. This setup induces the Markov chains $\bm{y}\leftrightarrow\bm{x}\leftrightarrow\bm{z}$ and $y_{i}\leftrightarrow\bm{x}\leftrightarrow\bm{z}$ for all $i=1,\dots,n$. We further assume that $\bm{z}$ can be decomposed into a set of variables, i.e., $\bm{z}=\{z_{j}\}_{j=1}^{m}$, with each $z_{j}$ also governed by a conditional distribution $p(z_{j}\mid\bm{x})$. From this point onward, we refer to the $z_{j}$ as the variables of $\bm{z}$.
@@
    Factors-Invariance: We say that a variable $z_{j}\in\bm{z}$ is factors-invariant with respect to a factor $y_{i}$ when it satisfies the following Markov chain:
    $\vskip-1.0pt\bm{\tilde{y}}_{i}\leftrightarrow y_{i}\leftrightarrow z_{j}\vskip-1.0pt$ (1)
    where $\bm{\tilde{y}}_{i}=\{y_{k}\}_{k\neq i}$ denotes the set of all other factors. This is equivalent to the conditional independence $I(z_{j};\bm{\tilde{y}}_{i}\mid y_{i})=0$, meaning that once $y_{i}$ is known, $z_{j}$ provides no additional information about the remaining factors $\bm{\tilde{y}}_{i}$. In other words, the distribution of $z_{j}$ is fully determined by the value of $y_{i}$ alone. This property is directly connected to modularity (Ridgeway & Mozer, 2018) (or disentanglement in Eastwood & Williams (2018)): a representation $z_{j}$ is said to be modular (or disentangled) if it captures at most one factor $y_{i}$. However, such definitions implicitly assume that the factors are independent. When dependencies exist among the factors—as is often the case in real-world scenarios—this definition becomes problematic. For instance, if two factors $y_{i}$ and $y_{k}$ are dependent, then $y_{k}$ inherently contains information about $y_{i}$. As a result, under the original definition, it would be impossible for $z_{j}$ to simultaneously be modular and fully informative about $y_{i}$, since it would necessarily encode some information about $y_{k}$ as well.
@@
    Nuisances-Invariance: We say that a variable $z_{j}\in\bm{z}$ is nuisances-invariant with respect to a factor $y_{i}$ and a set of nuisances $\bm{n}$ when it satisfies the following Markov chain:
    $\vskip-1.0pt\bm{n}\leftrightarrow y_{i}\leftrightarrow z_{j}\vskip-1.0pt$ (2)
    This is equivalent to the conditional independence $I(z_{j};\bm{n}\mid y_{i})=0$, meaning that once $y_{i}$ is known, $z_{j}$ remains unaffected by variations in the nuisances. This property is typically neither discussed nor measured in prior work—possibly due to the difficulty of estimating the distribution of $\bm{n}$. Despite this, it is an important criterion for disentangled representations. While factors-invariance ensures that $z_{j}$ is unaffected by other factors $y_{k}$ ($k\neq i$), we would also expect $z_{j}$ to be invariant to aspects of the input that are not factors of variation—i.e., the nuisances—if the representation is truly disentangled (Carbonneau et al., 2022).
@@
    Representations-Invariance: We say that a variable $z_{j}$ of the representation $\bm{z}$ is representations-invariant for a factor $y_{i}$ when it satisfies the following Markov chain:
    $\vskip-1.0pt\bm{\tilde{z}}_{j}\leftrightarrow z_{j}\leftrightarrow y_{i}\vskip-1.0pt$ (3)
    where $\bm{\tilde{z}}_{j}=\{z_{k}\}_{k\neq j}$ denotes all other variables in the representation. This is equivalent to the conditional independence $I(y_{i};\bm{\tilde{z}}_{j}\mid z_{j})=0$, which implies that $y_{i}$ can be entirely predicted from $z_{j}$ alone, without any additional information from the rest of the representation. This property is useful in downstream applications: for instance, if the goal is to predict $y_{i}$, one could use $z_{j}$ in isolation and ignore the remaining components of $\bm{z}$. Similarly, in controllable generative models, it would suffice to manipulate $z_{j}$ to control the value of $y_{i}$ in the generated output $\bm{x}$. This notion is closely related to compactness as defined in (Ridgeway & Mozer, 2018) (or completeness in Eastwood & Williams (2018)), where a representation $z_{j}$ is said to be compact (or complete) if it is the sole component encoding information about a factor $y_{i}$. As with factors-invariance, this property assumes independence among factors. In realistic scenarios where factors such as $y_{i}$ and $y_{k}$ are dependent, and represented by $z_{j}$ and $z_{l}$ respectively, it is natural for both $z_{j}$ and $z_{l}$ to carry information about both factors.
@@
    Explicitness We say that a representation $\bm{z}$ is explicit for a factor $y_{i}$ when it satisfies the following Markov chain:
    $\vskip-1.0pt\bm{x}\leftrightarrow\bm{z}\leftrightarrow y_{i}\vskip-1.0pt$ (4)
    This means that $I(y_{i};\bm{x}|\bm{z})=0$, i.e., $\bm{x}$ provides no information about $y_{i}$ when $\bm{z}$ is known or, equivalently, $\bm{z}$ contains all the information about $y_{i}$. This property is referred to as explicitness in Ridgeway & Mozer (2018) and as informativeness in Eastwood & Williams (2018). Unlike the other properties discussed, explicitness is not specific to disentangled representations—it is a desirable quality for representations more generally (Bengio et al., 2013).

### v3

@@ Abstract
Representation learning enables the discovery and extraction of underlying factors of variation from data. A representation is typically considered disentangled when it isolates these factors in a way that is interpretable to humans. Existing definitions and metrics for disentanglement often assume that the factors of variation are statistically independent. However, this assumption rarely holds in real-world settings, limiting the applicability of such definitions and metrics in real-world applications. In this work, we propose a novel definition of disentanglement grounded in information theory, which remains valid even when the factors are dependent. We show that this definition is equivalent to requiring the representation to consist of minimal and sufficient variables. Based on this formulation, we introduce a method to quantify the degree of disentanglement that remains effective in the presence of statistical dependencies among factors. Through a series of experiments, we demonstrate that our method reliably measures disentanglement in both independent and dependent settings, where existing approaches fail under the latter.
@@ 1 Introduction
In representation learning, data are commonly assumed to be characterized by a set of factors of variation (henceforth, factors) and nuisances. The distinction between them lies in task relevance: factors directly influence the outcome of interest, whereas nuisances do not. For instance, in a dataset of fruit images where the task is to describe the fruits, relevant factors may include fruit type, color, or size, while irrelevant nuisances might be the background color or the shadows cast. A large body of work has emphasized the importance of learning disentangled representations, where distinct dimensions of the representation correspond to different factors of variation (Schmidhuber, 1992; Bengio et al., 2013; Ridgeway, 2016; Lake et al., 2017; Achille and Soatto, 2018a). Such representations are valuable across a range of applications, including: (i) interpretability of learned features and model predictions (Hsu et al., 2017; Worrall et al., 2017; Gilpin et al., 2018; Zhang and Zhu, 2018; Liu et al., 2020; Zhu et al., 2021; Nauta et al., 2023), (ii) fairer decision-making by mitigating or removing biases in sensitive attributes such as gender or race (Calders and Žliobaitė, 2013; Barocas and Selbst, 2016; Kumar et al., 2018; Locatello et al., 2019a; Sarhan et al., 2020; Chen et al., 2023), and (iii) empowering generative models to synthesize data with controllable attributes (Yan et al., 2016; Paige et al., 2017; Huang et al., 2018; Dupont, 2018; Kazemi et al., 2019; Antoran and Miguel, 2019; Zhou et al., 2021; Shen et al., 2022; Almudévar et al., 2024).
The definition in the previous paragraph provides an intuitive, though imprecise, understanding of disentangled representations. It raises two fundamental questions: (i) what does it precisely mean to separate the factors, and (ii) how can we evaluate whether the factors are indeed separated? In Section 2, we review several works that attempt to address these two questions. Most of these approaches assume that the factors of variation are independent—both from one another and from the nuisances. Consequently, the corresponding definitions and evaluation metrics are grounded in this assumption of independence. Many of the benchmark datasets used to validate these methods also exhibit (approximately) independent factors (LeCun et al., 2004; Liu et al., 2015; Reed et al., 2015; Matthey et al., 2017; Burgess and Kim, 2018; Gondal et al., 2019), thereby making the independence-based definitions and metrics largely applicable in those contexts.
@@
  • We prove that these properties are directly connected to the concept of minimal and sufficient representations. Moreover, we argue that evaluating minimality and sufficiency provides a more meaningful assessment of disentanglement than analyzing the properties in isolation.
  • We propose a method to quantify the degree of minimality and sufficiency, enabling the evaluation of disentanglement in arbitrary representations.
Figure 1: If the factors are dependent, then—even with a perfectly disentangled encoder—each sub-representation will contain information about other factors. For example, if $z_{1}$ encodes the shape of a banana, the color is most likely yellow, possibly green if unripe, and rarely red. Likewise, if $z_{2}$ encodes the color yellow, the shape is most likely that of a banana or a lemon, but very unlikely that of a strawberry.
@@ 2 Related Work
Despite being a topic of great interest, there is no widely agreed-upon definition of disentangled representations. Intuitively, a disentangled representation separates the different factors of variation underlying the data (Desjardins et al., 2012; Bengio et al., 2013; Cohen and Welling, 2014; Kulkarni et al., 2015; Ridgeway, 2016). Early evaluations of disentanglement were largely based on visual inspection (Kingma and Welling, 2014). However, more precise definitions are necessary to understand the limitations and practical relevance of disentangled representations in downstream applications. One of the earliest and most widely cited definitions is given in Bengio et al. (2013), and has been adopted by several subsequent works (Higgins et al., 2017; Kim and Mnih, 2018; Kim et al., 2019; Suter et al., 2019). This definition states that a disentangled representation is one in which a change in a single dimension of the representation corresponds to a change in a single factor of variation, while remaining invariant to changes in other factors. An alternative view argues that disentanglement should be defined such that a change in a single factor of variation leads to a change in only one dimension of the representation (Kumar et al., 2018; Locatello et al., 2019b). More recent efforts define disentanglement based on a set of desirable properties. Notably, Ridgeway and Mozer (2018) and Eastwood and Williams (2018) independently proposed three key properties that a representation should satisfy to be considered disentangled. While the terminology differs slightly between the two works, the underlying ideas align: (i) Modularity(referred to as disentanglement in Eastwood and Williams (2018)): each representation variable captures at most one factor of variation; (ii) Compactness(completeness in Eastwood and Williams (2018)): each factor of variation is captured by a single representation variable; and (iii) Explicitness(informativeness in Eastwood and Williams (2018)): the representation retains all task-relevant information about the factors present in the input. All of the above definitions present two key limitations: (i) they do not account for invariance to nuisance factors that may be present in the input; and (ii) they assume that the factors of variation are independent of one another and of the nuisances—which generally does not hold in real-world data. We propose a set of properties that a representation must satisfy to be considered disentangled, explicitly addressing both of these aspects.
@@
  • Intervention-based Metrics. These methods construct data subsets that share a common factor of variation while differing in others. Representations for these subsets are computed and compared to produce a disentanglement score. Representative methods include $\beta$-VAE (Higgins et al., 2017), FactorVAE (Kim and Mnih, 2018), and R-FactorVAE (Kim et al., 2019). These metrics primarily assess modularity but do not account for compactness or explicitness (Sepliarskaia et al., 2019).
  • Information-based Metrics. These approaches estimate mutual information between the factors and individual variables in the representation. Widely used examples include the Mutual Information Gap (MIG) (Chen et al., 2018) and Robust MIG (Do and Tran, 2020). These metrics are mainly designed to capture compactness.
  • Predictor-based Metrics. These methods train regressors or classifiers to predict the factors from the learned representations. The resulting predictors are then analyzed to determine the contribution of each representation variable to each factor. Separated Attribute Predictability (SAP) (Kumar et al., 2018), focuses exclusively on compactness. Subsequently, more comprehensive frameworks have been proposed to evaluate multiple properties of disentanglement. Notably, Eastwood and Williams (2018) introduced DCI (Disentanglement, Completeness, and Informativeness), while Ridgeway and Mozer (2018) proposed an equivalent set: Modularity, Compactness, and Explicitness.
@@
The Information Bottleneck (IB) framework (Tishby et al., 1999) formalizes a trade-off between fidelity and compression in learned representations. A representation is deemed sufficient if it retains all information relevant to the target task, and minimal if it excludes all information irrelevant to it. Formally, given an input $X$, target variable $Y$, and representation $Z$, sufficiency corresponds to $I(Z;Y)=I(X;Y)$, while minimality requires $I(Z;X)=I(Z;Y)$. The IB principle has been leveraged in a range of works to design regularizers for representation learning (Alemi et al., 2017; 2018; Achille and Soatto, 2018b; a; Almudévar et al., 2025a), and has also been applied to promote disentanglement (Vera et al., 2018; Yamada et al., 2020; Jeon et al., 2021; Gao et al., 2021). To the best of our knowledge, however, this is the first work to formally define disentanglement in terms of minimal sufficient representations. In our formulation, each factor of variation is treated as a distinct target variable. Accordingly, we partition the overall representation vector into disjoint sets of dimensions, or ’sub-representations’. We define a representation as disentangled if each of these sub-representations is both minimal and sufficient with respect to its corresponding factor.
@@ 3 Desirable Properties of a Disentangled Representation
Notation.
We denote random variables with capital letters (e.g., $X$) and their realizations with lowercase letters (e.g., $x$). Vectors are denoted by bold lowercase letters (e.g., $\bm{x}$) and sets by calligraphic letters (e.g., $\mathcal{X}$). We use subscript indices to refer to specific components of a vector or set.
Problem Setup.
In our representation learning setup, we consider raw data $X\in\mathcal{X}$, which is fully explained by a set of task-relevant factors $Y=\{Y_{i}\}_{i=1}^{n}\in\mathcal{Y}$ and a set of nuisances $N\in\mathcal{N}$. The factors $Y$ correspond to the underlying sources or causes that influence the observed data and are relevant to the task at hand, whereas the nuisances $N$ refer to sources of variation in $X$ that are irrelevant to the task.
We define a representation $Z\in\mathcal{Z}$ as a random variable governed by the conditional distribution $p(\bm{z}\mid\bm{x})$. This setup induces the Markov chains $Y\leftrightarrow X\leftrightarrow Z$ and $Y_{i}\leftrightarrow X\leftrightarrow Z$ for all $i=1,\dots,n$. We further assume that $Z$ can be decomposed into a set of $m$ variables, $Z=\{Z_{j}\}_{j=1}^{m}$, where each $Z_{j}$ is governed by the conditional distribution $p(z_{j}\mid\bm{x})$.
Desirable Properties
@@
    Factors-Invariance: We say that a variable $Z_{j}\in Z$ is factors-invariant with respect to a factor $Y_{i}$ when it satisfies the following Markov chain:
    $\vskip-1.0pt\tilde{Y}_{i}\leftrightarrow Y_{i}\leftrightarrow Z_{j}\vskip-1.0pt$ (1)
    where $\tilde{Y}_{i}=\{Y_{k}\}_{k\neq i}$ denotes the set of all other factors. This is equivalent to the conditional independence $I(Z_{j};\tilde{Y}_{i}\mid Y_{i})=0$, meaning that once $Y_{i}$ is known, $Z_{j}$ provides no additional information about the remaining factors $\tilde{Y}_{i}$. In other words, the distribution of $Z_{j}$ is fully determined by the value of $Y_{i}$ alone. This property is directly connected to modularity (Ridgeway and Mozer, 2018) (or disentanglement in Eastwood and Williams (2018)): a representation $Z_{j}$ is said to be modular (or disentangled) if it captures at most one factor $Y_{i}$. However, such definitions implicitly assume that the factors are independent. When dependencies exist among the factors—as is often the case in real-world scenarios—this definition becomes problematic. For instance, if two factors $Y_{i}$ and $Y_{k}$ are dependent, then $Y_{k}$ inherently contains information about $Y_{i}$. As a result, under the original definition, it would be impossible for $Z_{j}$ to simultaneously be modular and fully informative about $Y_{i}$, since it would necessarily encode some information about $Y_{k}$ as well.
@@
    Nuisances-Invariance: We say that a variable $Z_{j}\in Z$ is nuisances-invariant with respect to a factor $Y_{i}$ and a set of nuisances $N$ when it satisfies the following Markov chain:
    $\vskip-1.0ptN\leftrightarrow Y_{i}\leftrightarrow Z_{j}\vskip-1.0pt$ (2)
    This is equivalent to the conditional independence $I(Z_{j};N\mid Y_{i})=0$, meaning that once $Y_{i}$ is known, $Z_{j}$ remains unaffected by variations in the nuisances. This property is typically neither discussed nor measured in prior work—possibly due to the difficulty of estimating the distribution of $N$. Despite this, it is an important criterion for disentangled representations. While factors-invariance ensures that $Z_{j}$ is unaffected by other factors $Y_{k}$ ($k\neq i$), we would also expect $Z_{j}$ to be invariant to aspects of the input that are not factors of variation—i.e., the nuisances—if the representation is truly disentangled (Carbonneau et al., 2022).
@@
    Representations-Invariance: We say that a variable $Z_{j}$ of the representation $Z$ is representations-invariant for a factor $Y_{i}$ when it satisfies the following Markov chain:
    $\vskip-1.0pt\tilde{Z}_{j}\leftrightarrow Z_{j}\leftrightarrow Y_{i}\vskip-1.0pt$ (3)
    where $\tilde{Z}_{j}=\{Z_{k}\}_{k\neq j}$ denotes all other variables in the representation. This is equivalent to the conditional independence $I(Y_{i};\tilde{Z}_{j}\mid Z_{j})=0$, which implies that $Y_{i}$ can be entirely predicted from $Z_{j}$ alone, without any additional information from the rest of the representation. This property is useful in downstream applications: for instance, if the goal is to predict $Y_{i}$, one could use $Z_{j}$ in isolation and ignore the remaining components of $Z$. Similarly, in controllable generative models, it would suffice to manipulate $Z_{j}$ to control the value of $Y_{i}$ in the generated output $X$. This notion is closely related to compactness as defined in (Ridgeway and Mozer, 2018) (or completeness in Eastwood and Williams (2018)), where a representation $Z_{j}$ is said to be compact (or complete) if it is the sole component encoding information about a factor $Y_{i}$. As with factors-invariance, this property assumes independence among factors. In realistic scenarios where factors such as $Y_{i}$ and $Y_{k}$ are dependent, and represented by $Z_{j}$ and $Z_{l}$ respectively, it is natural for both $Z_{j}$ and $Z_{l}$ to carry information about both factors.
@@
    Explicitness We say that a representation $Z$ is explicit for a factor $Y_{i}$ when it satisfies the following Markov chain:
    $\vskip-1.0ptX\leftrightarrow Z\leftrightarrow Y_{i}\vskip-1.0pt$ (4)
    This means that $I(Y_{i};X|Z)=0$, i.e., $X$ provides no information about $Y_{i}$ when $Z$ is known or, equivalently, $Z$ contains all the information about $Y_{i}$. This property is referred to as explicitness in Ridgeway and Mozer (2018) and as informativeness in Eastwood and Williams (2018). Unlike the other properties discussed, explicitness is not specific to disentangled representations—it is a desirable quality for representations more generally (Bengio et al., 2013).